OpenStudy (anonymous):

Write an equation of the line, in the slope-intercept form (y=mx+b) with a slope of 4 and passes through point (3,6)

hero (hero):

The slope intercept form of a line is \(y = mx + b\) To write the equation of a specific line we need to input the values of \(m\) and \(b\) In this case, we have \(m\) but we need \(b\). Luckily, we've been given a point that is on the line, so we will input the values for the point and the slope, then solve for b: \(6 = 4(3) + b\) \(6 = 12 + b\) \(6 - 12 = b\) \(-6 = b\) Now that we have the values for \(m\) and \(b\) we can write the equation of the line: \(y = 4x - 6\)

OpenStudy (anonymous):

the equation of a line is given by (y-y0)=m(x-x0) where x0=3 and y0=6 this gives us: y-6=4(x-3) which simplifies to: y=4x-6